pls solve this question...its urgent.....The rate of one knot equals one nautical mile per hour. One nautical mile is 1852 meters. What is the speed in meters per second of a ship traveling at 20 knots?

Answer by mananth(16949)

1 knot = 1852 meters / hour
20 knots = 1852*20 = 37040 meters / hour
37040 /3600 = 10.33 meters / second